A student determines the chromium(II) content of a solution by first precipitating it as chromium(II) hydroxide, and then decomposing the hydroxide to chromium(II) oxide by heatingHow grams of chromium()oxide should the student obtain if her solution contains 55.0 of 0.569 M chromium(II) nitrate?


1
Expert's answer
2021-05-06T07:37:08-0400

Moles ="\frac {36×0.507}{1000}=\frac {Mass of CrO} {67.996g/mol}"


Mass ="\frac {36×0.507×67.996}{1000}"


Mass =1.241g of CrO obtained.